OpenStudy (anonymous):

Use a graph to solve the equation on the interval [−2π, 2π]. sec(x)= sqrt (2)

OpenStudy (aum):

Graph \(\Large \sec(x) - \sqrt{2}\) and see where it crosses the x-axis in the interval \(\Large[-2\pi, 2\pi]\).